hydration energy of Na2SO4 is more than its lattice energy

Grade:12

1 Answers

ruchi yadav
10 years ago
In Na2SO4, q1 × q2 =1 × 2. While in BaSO4, q1 × q2 =2 × 2 = 4. Since lattice energy ? q1q2, lattice
energy of BaSO4 is about twice that of Na2SO4. In Na2SO4, lattice energy is less than hydration energy, hence soluble in water.


Thank You
Ruchi
Askiitians Faculty